OpenStudy (anonymous):

how much pure acid should be mixed with 2 gallons of an 80% acid solution in order to get a 90% acid solution? to get a 90% acid solution you need how many gallons of pure acid

OpenStudy (anonymous):

i may be wrong,because i studied way back. acid in 2 gallons of solution\[=\frac{ 80 }{ 100 } \times 2=1.60 gallons\] let x be the number of pure acid added to make it 90%. \[\frac{ 1.60+x }{ 2+x }=\frac{ 90 }{ 100 }\] 160+100x=180+90x 100x-90x=180-160 10x=20 x=20/10=2 Hence 2 gallons of pure acid should be added to 2 gallons of 80 % solution.